### v1

# Can Large Language Models Be Trusted as Black-Box Evolutionary Optimizers for Combinatorial Problems?

Abstract
Evolutionary computation excels in complex optimization but demands deep domain knowledge, restricting its accessibility. Large Language Models (LLMs) offer a game-changing solution with their extensive knowledge and could democratize the optimization paradigm. Although LLMs possess significant capabilities, they may not be universally effective, particularly since evolutionary optimization encompasses multiple stages. It is therefore imperative to evaluate the suitability of LLMs as evolutionary optimizer (EVO). Thus, we establish a series of rigid standards to thoroughly examine the fidelity of LLM-based EVO output in different stages of evolutionary optimization and then introduce a robust error-correction mechanism to mitigate the output uncertainty. Furthermore, we explore a cost-efficient method that directly operates on entire populations with excellent effectiveness in contrast to individual-level optimization. Through extensive experiments, we rigorously validate the performance of LLMs as operators targeted for combinatorial problems. Our findings provide critical insights and valuable observations, advancing the understanding and application of LLM-based optimization.
Keywords Evolutionary optimization  $\cdot$ Large language models  $\cdot$ Combinatorial problems
1 Introduction
The advancement of computing technology, especially large language models (LLMs) [1, 2], has made it possible to revolutionize traditional programming methodologies [3, 4, 5] and reduce human labor. This vision is already materializing in the field of optimization [6], where optimization of combinatorial problems is achieved not through mathematical analysis or traditional programming but via prompt engineering only. Evolutionary optimization has proven to be an exceptionally powerful technique for tackling complex problems due to its simplicity and effectiveness [7, 8]. Despite their benefits, implementing evolutionary optimization is labor-intensive and demands deep domain expertise for precise execution [9]. This challenge intensifies with constrained problems or when tailoring methods for specific needs. Such difficulty is especially pronounced for non-technical users, presenting a major barrier to accessibility and effective application. Therefore, we ask the critical question of whether LLMs can directly serve as evolutionary optimizer (EVO).
To date, some preliminary studies have explored the use of LLMs as crossover operators [10], mutation operators [11], and in combination [12]. Despite these advancements [13, 14], LLMs have shown variable performance across different domains; they excel in some areas [15, 16, 17] but not in tasks requiring precise arithmetic and logical reasoning [18, 19]. For most researchers, it is challenging to train a large model from scratch targeted to a specific problem or even fine-tune it due to the huge computational demands. Thus, when deploying these models for specific tasks, it is imperative to conduct a thorough analysis to ascertain whether LLMs are genuinely suitable for the task at hand.
While these efforts have opened up LLMs as the black-box EVO, there is still a lack of comprehensive investigation to evaluate the effectiveness and reliability of LLMs as EVO, and to identify the factors affecting their performance. Huang et al. [20] conducted an investigation on LLMs as normal optimizers on different problems, but it is only targeted at one-shot optimization, and some important aspects such as reliability, scalability, and computational cost remain underexplored. Thus, our work here aims to serve as a thorough assessment of the performance of LLMs as operators in all stages of evolutionary optimization, some of which, like LLM-based initialization and selection, still remain underexplored in the existing literature. We aim to provide insights into the suitability of LLMs, clarifying their capabilities and limitations. Our evaluations compare LLM-based approaches with software-based evolutionary computation to assess the accuracy of LLMs in meeting expected outcomes and their effectiveness in fulfilling customized requirements.
As evolution optimization is an iterative process, an error in any step will yield a cascading effect, leading to a failure of optimization. Recognizing the inherently probabilistic nature of LLMs outputs [21, 22, 23] and the high requirements on the quality of the generated solutions during evolutionary optimization, we develop different stringent sets of standards for LLMs outputs in various levels to rigorously measure solutions in terms of their format, diversity, and conformity to problem constraints. We also introduce a set of corresponding error repair mechanisms with precisely tailored prompts to enhance the reliability of the LLM-based EVO.
In addition, we explore a cost-effective method that utilises the entire population as optimization unit for the LLM-based reproduction operator and compare this method with another common approach that implements optimization on individual-by-individual. We specifically analyze the quality of generated solutions by these two distinct ways as well as their computational overhead. Some of our findings in this work are summarized as follows:
$\bullet$ LLMs can perform the selection, crossover, and mutation operations well. Their performance is relatively independent of dataset size and type, being more influenced by hyperparameters such as population size and solution size instead.
$\bullet$ LLMs may not be suitable for tasks involving intensive computation, such as initialization, and their performance tends to degrade greatly as the dataset size increases.
$\bullet$ The performance of LLMs relies on the amount of input data, and LLMs may encounter various issues during the evolutionary optimization process, making correction mechanisms indispensable for successful optimization.
2 LLM-based evolutionary optimizer
As in the combinatorial problem within the evolutionary optimization framework, the solution representation is generally defined as
$[\text{X}_{1},\text{X}_{2},...,\text{X}_{n}],$ (1)
where X depends on the specific task and $n$ is the solution size [24, 25]. In this work, we choose a node selection problem [26, 27] as the illustrative example. The problem is defined as: Given a graph $G=(V,E)$, where $V$ represents the set of nodes and $E$ represents the set of edges, the objective is to find a subset of nodes $S\subseteq V$ that maximizes the influence across the network. As for network-related problems, the solution representation can simply be
$[\text{Node}_{1},\text{Node}_{2},...,\text{Node}_{n}].$ (2)
Note that our focus is not on solving any specific network-related problem but on evaluating the performance of LLMs as operators for combinatorial problems. The reason for selecting this problem is that it is general enough to provide insights and observations for LLMs as EVO due to the representation similarity on combinatorial problems. To enhance generality, ‘element’ will replace ‘node’, and ‘dataset’ will replace ‘graph’ in the rest discussion. This generality is also reflected in our prompt engineering where little network domain knowledge is needed (please refer to the Appendix). For the selection, the inputs to LLMs is
$\{([\text{X}_{1}^{(1)},\ldots,\text{X}_{n}^{(1)}],f_{1}),\ldots,([\text{X}_{1}^{(k)},\ldots,\text{X}_{n}^{(k)}],f_{k})\},$ (3)
where $k$ is the population size and $\text{X}_{i}^{(j)}$ refers to the $i$-th element in the $j$-th solution in the population. It can be simplified as
$\{(S_{1},f_{1}),...,(S_{k},f_{k})\},$ (4)
where $S_{i}$ is the index of solution $[\text{X}_{i}^{(1)},\ldots,\text{X}_{n}^{(i)}]$ and $f_{i}$ is its fitness value.
Let $\mathbf{F}$ be combination of LLM-based evolutionary operators, i.e.,
$\mathbf{F}=\{\mathbf{F}_{I},\mathbf{F}_{S},\mathbf{F}_{C},\mathbf{F}_{M}\},$ (5)
where $\mathbf{F}_{I}$, $\mathbf{F}_{S}$, $\mathbf{F}_{C}$ and $\mathbf{F}_{M}$ indicate LLM-based initialization, selection, crossover, and mutation, respectively.
Figure 1: (A) The diagram of LLM-based EVO with the proposed validation and repair mechanism. All four phases in evolution optimization along with repair are based on LLMs. (B) The illustration of population- and individual-level LLM-based EVO. (C) An example of errors encountered in population-level LLM-based crossover. The error message (if any), customized corrected prompt, and previous deficient output will be provided to LLMs for repair.
The first phase, Initialization is defined as $\mathcal{P}_{0}=\mathbf{F}_{I}(G)$, where $\mathcal{P}_{0}$ refers to the initialized population and $G$ refers to the input data. A strategic reduction in search space can effectively speed up convergence and improve effectiveness [28]. However, based on current research [18], it is not feasible to feed datasets directly to LLMs, either in the form of adjacency matrices or text descriptions, because the reasoning ability of current LLMs on graphs remains limited. Therefore, instead of providing LLMs with just a dataset, we opt to input element IDs along with easily accessible metrics. Following this, LLMs will be instructed to (1) rank the element based on a specific metric; (2) select a percentage of top elements as candidates; and (3) sample filtered elements to form the initialized population. Several basic restrictions must be followed to ensure the quality of the initialized population: (1) Each solution in the population should contain the same number of elements with no duplicates; (2) The solutions within the population should be as diverse as possible; and (3) elements outside the candidate set are not allowed to appear in the population.
The Selection phase in evolutionary optimization is a crucial step where individuals from the current population are selected as the next generation. This is represented as $\mathcal{P}_{l+1}=\mathbf{F}_{S}(\mathcal{P}_{l},f(\mathcal{P}_{l}))$, where $\mathcal{P}_{l+1}$ refers to the population after applying operators of $\{\mathbf{F}_{C},\mathbf{F}_{M}\}$ on $\mathcal{P}_{l}$ and $f(\mathcal{P}_{l})$ represents the fitness values of solutions in $\mathcal{P}_{l}$. Extensive studies on selection strategies, such as roulette wheel selection, aim to preferentially choose individuals with higher fitness scores. This operation allows the population to evolve towards an optimal solution over successive generations. In $\mathbf{F}_{S}$, we will not rely on any specified strategy but instead follow the common principles: (1) Solutions with low fitness are not allowed to be selected; and (2) Each solution can be selected multiple times but not excessively, to maintain diversity.
As for reproduction, the individual-level LLM-based operators of Crossover $\mathbf{F}_{C}^{S}$ and Mutation $\mathbf{F}_{M}^{S}$ are
$\begin{split}&\mathcal{P}_{l+1}=\bigcup_{\mathcal{S}_{i},\mathcal{S}_{j}\in\mathcal{P}_{l}}\mathbf{F}_{C}^{S}(\mathcal{S}_{i},\mathcal{S}_{j}),\\ &\mathcal{P}_{l+1}=\bigcup_{\mathcal{S}\in\mathcal{P}_{l}}\mathbf{F}_{M}^{S}(\mathcal{S}),\end{split}$ (6)
where $\mathcal{S}$ denotes the solution in population $\mathcal{P}$. The individual-level operator is the conventional method for implementing evolutionary optimization [12, 13]. While this individual-level method provides a precise and controlled way to reproduce the population, it may suffer from two problems: (i) handling solutions individually or in pairs might not scale efficiently as the population size grows in the framework of LLM-based EVO; and (ii) LLMs may lack broader contextual information about the population’s overall diversity, which could impede their ability to optimize effectively. To resolve these critical issues, we propose here a new population-level optimization, defined as follows:
$\begin{split}&\mathcal{P}_{l+1}=\mathbf{F}_{C}^{P}(\mathcal{P}_{l}),\\ &\mathcal{P}_{l+1}=\mathbf{F}_{M}^{P}(\mathcal{P}_{l}).\\ \end{split}$ (7)
Here we distinguish between these two methods: All evolutionary optimizations rely on populations, typically selecting one or a pair of individuals from the population iteratively to input into LLMs, which is referred to as individual-level in our work. In contrast, our proposed population-level method is to input the entire population into LLMs for optimization.
One unique advantage of the proposed population-level LLM-based optimization lies in its global awareness of the diversity of the entire population. This enables the potential for intelligent reproduction, specifically selecting individuals most likely to produce optimal solutions for recombination, rather than relying on random selection or investing substantial effort in designing complex algorithms to choose individuals, as is typical in individual-level reproduction, whether based on code or LLMs.
2.1 Computational cost of population- and individual-level LLM-based EVO
As we aim to assess the fidelity of LLM-based EVO, the benchmark is based on precise coding methods. It is not applicable to compare their computational costs and efficiency since LLMs rely on online resources (e.g., token count) while code-based methods depend on local hardware. Then, we will discuss the difference between population- and individual-level optimization regarding computational cost.
[…]
Mutation: As for individual-level mutation, it will run $\mathbf{F}_{M}^{S}$ for $\mathcal{N}_{p}$ times since it will mutate the solution individual by individual. Thus, its cost difference with population-level mutation is $\mathcal{V}(\mathcal{T}_{M}^{S}*\mathcal{N}_{p})-\mathcal{V}(\mathcal{T}_{M}^{P})\approx(\mathcal{N}_{p}-1)*\mathcal{V}(\mathcal{T}_{M}^{S})$. Clearly, the proposed population-level method can save roughly $\frac{\mathcal{N}_{p}-2}{2}*\mathcal{V}(\mathcal{T}_{C}^{S})+(\mathcal{N}_{p}-1)*\mathcal{V}(\mathcal{T}_{M}^{S})$ in each round as compared to individual-level method, and this advantage in cost will be further amplified when population size and evolution round are set to a high number.
Notably, there are several restrictions that must be followed during reproduction: (1) The new solution should not contain duplicate/invalid elements; (2) The new solution should maintain the same size as the previous solution; and (3) The population size should remain the same as the previous population (only applicable to the population-level EVO).
Error Type Error Index Error Description
Format Error E1 The output is not in the required format.
E2 The output contains non-integer elements.
Critical Error E3 The selected candidates significantly deviate from the ground truth.
E4 The size of candidates falls significantly short of meeting the requirements.
E5 The size of the population falls significantly short of meeting the requirements.
E6 The selected population contains one solution too many times.
E7 Any solution appears in the population where all elements are the same.
E8 The number of different elements in the solution changes significantly.
E9 The number of different solutions in the population changes significantly.
Moderate Error E10 The size of some solutions fails to meet the requirement.
E11 The solution (population) after the operation remains the same as before.
E12 The size of the population fails to meet the requirement.
E13 The new solution contains duplicated elements.
E14 The new solution contains invalid elements not found in candidate nodes.
E15 The selected population contains those with very low fitness.
Table 1: The errors of LLM outputs during evolutionary optimization.
2.2 Error repair
Phase Format Error Critical Error Moderate Error
E1 E2 E3 E4 E5 E6 E7 E8 E9 E10 E11 E12 E13 E14 E15
Candidate selection - - - - - - - - - - -
Initialization - - - - - - - -
Selection - - - - - - - - - -
Crossover (P) - - - - -
Crossover (S) - - - - - - - -
Mutation (P) - - - -
Mutation (S) - - - - - - -
Table 2: The required validations for the LLM output in different phases of evolutionary optimization.
The probabilistic nature of LLMs can lead to occasional undesirable results. As evolutionary optimization is an iterative process, an error early in the process can ‘blow up’ in subsequent cycles, creating a cascading effect to ruin the optimization process. We therefore investigate the possible errors and set rigid standards to validate the LLM outputs in different phases. These errors are categorized as format errors and quality errors, with the latter further divided into critical errors and moderate errors. Specifically,
  • Format Error refers to generated solutions that fail to adhere to required structures.
  • Critical Error refers to solutions that, while structurally correct, may cause system failures, demanding immediate resolution to prevent extensive damages.
  • Moderate Error refers to solutions that can cause minor disruptions but generally will not lead to complete failure of optimization.
The details of the different types of errors are listed in Table 1 (Please refer to the Appendix for more explanation) and an example of possible errors encountered in the population-level crossover is given in Figure 1(C). The outputs generated in different phases will undergo different validations due to the varying requirements for each phase. In addition, the output requirements are different when the input is an individual solution (denoted as $S$) and the entire population (denoted as $P$). The detail is shown in Table 2.
During optimization, particularly in environments with numerous constraints, managing invalid solutions is critical in ensuring successful optimization. To this end, we introduce a robust repair mechanism that maintains solution feasibility and ensures that each iteration contributes positively toward finding an optimal solution. Let $\mathcal{O}$ be the output from $\{\mathbf{F}_{I},\mathbf{F}_{S},\mathbf{F}_{C},\mathbf{F}_{M}\}$, the refined output is defined as
$\mathcal{O}\leftarrow\mathbf{F}_{R}(\mathcal{O},E_{x},R_{x}),$ (8)
where $\mathbf{F}_{R}$ refers to the LLM-based repair operator. $E_{x}$ and $R_{x}$ denote the error message and targeted repair prompt.
During the optimization, each output undergoes the three aforementioned examinations. If any error is detected, the repair mechanism is triggered. We will check and repair each type of error individually. Format errors and critical errors are particularly detrimental. The former interrupts the entire optimization process, and the latter can heavily corrupt the population, diverting the optimization from its optimal path. We will avoid repairing these two types of errors due to the complexity involved. Instead, a new generation is directly requested to obtain a valid solution. Due to their destructive impact, solutions failing these checks cannot be used and the previous phase’s solution will be used for the next phase.
In contrast, solutions with moderate errors that are not successfully repaired are allowed to enter into the next phase. This approach is adopted because (1) the error can always be mitigated after repair; and (2) moderate errors do not yield cascading effects on the entire optimization, and there will still be an opportunity to repair them in the next phases.
2.3 Solution evaluation
After the output undergoes the check and, if necessary, repair process, we categorize it into the following cases:
  • Approved ($\mathsf{Q}_{\mathsf{app}}$): The solution meets all requirements and standards perfectly, with no errors or deficiencies. It is ready for implementation without any modifications.
  • Repaired ($\mathsf{Q}_{\mathsf{rep}}$): The solution had minor issues that did not meet the necessary standards, but these have been addressed, and it now meets the required criteria.
  • Acceptable ($\mathsf{Q}_{\mathsf{acc}}$): The solution contains flaws that fail to be completely corrected, but it still functions adequately and meets the minimum necessary criteria for use, albeit not optimally.
  • Rejected ($\mathsf{Q}_{\mathsf{rej}}$): The solution has format or critical error even though after an attempt of repair, rendering it completely unusable.
For the format and critical errors, the acceptable case $\mathsf{Q}_{\mathsf{acc}}$ is not applicable due to their destructive impact on the population. If a solution encounters these errors and cannot be repaired, it must be rejected. In this case, the output with moderate error from last check will pass through to the next phase if repair fails.
3 Experimental studies
The fidelity and reliability of LLM-based EVO of different stages are validated on various datasets in different settings. The population size and the evolution round are set to 30. The number of tests for candidate selection is 100, and the results of evolutionary optimization are averaged from 10 independent simulations. The structural information of the tested dataset is listed in Table 3. More details of the parameters are given in the Appendix.
Dataset Dolphins Netscience Erods Email Astro
$\boldsymbol{|V|}$ 62 379 433 1,005 14,845
$\boldsymbol{|E|}$ 159 914 1,314 25,571 119,652
Table 3: The structural information of network dataset.
3.1 Optimization validation of LLM-based EVO
3.1.1 Initialization:
Table 4 shows the results of pass ratio of format and critical check in the candidate selection, consisting of ranking and filtering. As seen, LLMs effective on small datasets fail to generate valid outputs for large datasets. For example, on Dolphins, LLMs achieve 100.0% $\mathsf{Q}_{\mathsf{acc}}$ on both checks. We compare it with larger networks and observe an obvious decrease in pass ratios; particularly, the critical check pass drops to about 40%. Thus, it can be seen that LLMs’ performance in calculation tasks, like candidate selection, will decline as the amount of input data increases.
Dataset Dolphins Netscience Erods
Format check pass 100.0% 95.0% 94.0%
Critical check pass 100.0% 41.0% 42.0%
Table 4: The validation of LLM-based candidate selection.
The LLM-based initialization shows superior fitness at the starting point and during the optimization compared to random initialization, as shown in Figure 2. This result shows that LLMs can effectively handle population initialization, reducing human effort for networks of hundreds of nodes.
Figure 2: The optimization of different initialization methods.
3.1.2 Selection:
Figure 3 compares the selection performance of LLM-based selection with other selection strategies. The LLM-based selection demonstrates the highest fitness across the entire optimization generation, with only the tournament strategy achieving comparable performance. It can be deduced that LLMs is effective in decision-making tasks such as in the selection phase.
Figure 3: The optimization of different selection methods.
3.1.3 Reproduction:
For fairness, we enhanced the baseline code-based EVO shown in Figure 4 by ensuring it does not violate constraints. As observed, the LLM-based EVO can achieve nearly the same performance as software-based evolutionary optimizers, reflected by their similar fitness. Additionally, the choice of model significantly impacts performance, as evidenced by the poor results of the population-level LLM-based EVO by GPT-3.5.
Figure 4: The comparison of different optimizers.
Figure 5: The validation of moderate error in different phases.
3.2 Reliability analysis
Netscience Format Error (GPT-4.0) Critical Error (GPT-4.0) Format Error (GPT-3.5) Critical Error (GPT-3.5)
$\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$
Initialization 100.0% 0.0% 0.0% 97.0% 2.0% 1.0% 98.0% 2.0% 0.0% 84.0% 16.0% 0.0%
Selection 100.0% 0.0% 0.0% 99.3% 0.7% 0.0% 100.0% 0.0% 0.0% 95.0% 4.0% 1.0%
Crossover (P) 99.6% 0.0% 0.4% 97.6% 2.3% 0.0% 95.6% 4.4% 0.0% 78.0% 13.6% 8.4%
Mutation (P) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 98.0% 2.0% 0.0% 92.3% 3.3% 4.4%
Crossover (S) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 99.6% 0.4% 0.0% 100.0% 0.0% 0.0%
Mutation (S) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 96.0% 3.7% 0.3% 99.9% 0.0% 0.1%
Table 5: The validations of format and critical errors for the LLMs output in different phases.
Table 5 shows that the ratio of $\mathsf{Q}_{\mathsf{app}}$ regarding format error is consistently above 95% in all phases. Meeting the requirement of critical errors is more challenging as more criteria are involved. GPT-4.0 performs exceptionally well, achieving a $\mathsf{Q}_{\mathsf{rej}}$ below 1%. Even when issues are occasionally detected, they can be repaired, as seen in the initialization case. In contrast, LLM-based EVO enhanced by GPT-3.5 has a relatively low $\mathsf{Q}_{\mathsf{acc}}$ and has nearly 10% $\mathsf{Q}_{\mathsf{rej}}$ in critical error during the population-level crossover. Figure 5 also exhibits the similar result as format and critical check. The ratio of $\mathsf{Q}_{\mathsf{acc}}$ in moderate error of GPT-4.0 is higher than that obtained by GPT-3.5. It can be concluded that the quality of LLMs output is promising but strongly dependent on model capability. In phases that demand processing large data amount, such as population-level crossover and mutation, the LLM-based operator only achieves a maximum accuracy ($\mathsf{Q}_{\mathsf{acc}}$) of 92.3%. However, in phases requiring less data processing, it can maintain a minimum accuracy of 95%. There are also fewer errors for individual-level EVO than population-level EVO. We can conclude that the reliability of LLMs is mainly affected by the scale of the processed data.
Operation Dataset Format Error (GPT-4.0) Critical Error (GPT-4.0) Moderate Error (GPT-4.0)
$\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{acc}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$
Crossover Erods 99.3% 0.3% 0.4% 98.6% 1.4% 0.0% 50.0% 30.1% 19.3% 0.5%
Email 100.0% 0.0% 0.0% 99.3% 0.7% 0.0% 57.2% 24.0% 17.9% 0.8%
Astro 99.3% 0.7% 0.0% 98.3% 1.7% 0.0% 55.8% 26.2% 18.1% 0.8%
Mutation Erods 100.0% 0.0% 0.0% 99.3% 0.7% 0.0% 80.8% 14.6% 4.3% 0.1%
Email 98.0% 2.0% 0.0% 94.3% 4.3% 1.4% 82.1% 14.3% 3.2% 0.3%
Astro 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 83.3% 14.0% 2.7% 0.1%
Table 6: The validations of format and critical errors for the LLMs output during crossover and mutation.
3.3 Scalability analysis
The scalability of LLM-based operators is subjected to the input data, which varies across different stages.
Initialization: As illustrated in Table 4, the effectiveness of candidate selection is significantly influenced by the size of the dataset. The input scale for LLMs during initialization (both candidate selection and sampling) increases linearly with the dataset size. For example, given a network of 10,000 nodes, we must input all their IDs with any specific metrics to LLMs for ranking, filtering, and sampling. It is not practical to input such large amounts of numerical data into LLMS and have them do such complex operations, not to mention the huge token overhead, meaning that LLMs may not be suitable for this data-intensive task.
Previous experiments indicate that LLMs are highly sensitive to the amount of input data. However, apart from the initialization phase, the input to LLMs is not the dataset itself but the population. Thus, the performance of LLM-based selection and reproduction relies more on hyperparameters than on the scale or type of dataset.
Selection: No matter what the dataset is, the input to LLM is always $\{(S_{1},f_{1}),...,(S_{k},f_{k})\}$. Thus, the relevant factor to LLM-based selection is the population size. A larger population size increases the amount of data LLMs need to process, but the common setting of this parameter is manageable to LLMs [29, 24]. Therefore, LLM-based selection has great applicability in this decision-making task given the promising result in Figure 3.
Crossover and Mutation: For these two phases, the input to LLMs is either $[\text{X}_{1},\text{X}_{2},...,\text{X}_{n}]$ (individual-level) or $\{[\text{X}_{1}^{(1)},\ldots,\text{X}_{n}^{(1)}],\ldots,[\text{X}_{1}^{(k)},\ldots,\text{X}_{n}^{(k)}]\}$ (population-level). As such, the input data amount, the main factor affecting LLMs’ performance, is dependent on the solution size and population size (only applicable to population-level optimization). The population size and solution size are empirical and will unnecessarily increase with the increase in the scale of datasets. Thus, we can conclude that the dataset scale and the LLM’s performance are not directly related, which is demonstrated in Table 6 where the reliability of LLM-based EVO in the large dataset Astro is comparable to that in the two smaller networks.
Evidence of LLMs sensitivity to hyperparameters can be found in Table 7, where the population size is set to $\mathcal{P}_{1}=30$ and $\mathcal{P}_{2}=10$, and the solution size is set to $\mathcal{S}_{1}=10$ and $\mathcal{S}_{2}=5$. Reducing either the population size or the size of a single solution helps to minimize errors, a trend consistent for both crossover and mutation phases.
One potential factor affecting performance as dataset size increases is the representation of element IDs in the solution. For example, given two datasets with element counts of up to $10^{3}$ and $10^{4}$, the maximum number of digits required to represent element IDs is 4 and 5, respectively. It means that the growth in element ID digits increases logarithmically rather than proportionally with dataset size. Therefore, we can conclude the scalability of LLM-based EVO in the reproduction phase is excellent, which is also demonstrated in Table 6, where there is no obvious difference between Astro and other smaller datasets regarding three types of check.
Netscience Format Error (GPT-4.0) Critical Error (GPT-4.0) Moderate Error (GPT-4.0)
$\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{acc}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$
C-$\boldsymbol{(\mathcal{P}_{1},\mathcal{S}_{1})}$ 96.7% 0.0% 3.3% 97.6% 2.4% 0.0% 51.8% 27.0% 20.4% 0.8%
C-$\boldsymbol{(\mathcal{P}_{1},\mathcal{S}_{2})}$ 100.0% 0.0% 0.0% 99.4% 0.6% 0.0% 73.3% 18.9% 7.6% 0.2%
C-$\boldsymbol{(\mathcal{P}_{2},\mathcal{S}_{1})}$ 99.3% 0.7% 0.0% 99.3% 0.7% 0.0% 93.2% 4.8% 1.5% 0.5%
M-$\boldsymbol{(\mathcal{P}_{1},\mathcal{S}_{1})}$ 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 77.7% 18.4% 3.8% 0.1%
M-$\boldsymbol{(\mathcal{P}_{1},\mathcal{S}_{2})}$ 100.0% 0.0% 0.0% 99.6% 0.4% 0.0% 92.9% 6.3% 0.8% 0.0%
M-$\boldsymbol{(\mathcal{P}_{2},\mathcal{S}_{1})}$ 100.0% 0.0% 0.0% 98.0% 2.0% 0.0% 95.7% 3.8% 0.5% 0.0%
Table 7: The validations for different errors for the LLMs output with different settings during crossover and mutation.
3.4 Ablation study
Figure 6 presents the ablation results of implementing a repair mechanism on the population-level LLM-based EVO across three datasets. We only test the moderate error as format error and critical error are so severe that the optimization cannot proceed as normal, and lead to consistently poor outcomes. For both datasets, the inclusion of a repair mechanism consistently outperforms the absence of one, as indicated by the higher average fitness achieved across generations. The observed stagnation in the scenarios without a repair mechanism indicates the impact of errors on optimization and suggests the importance of repair procedures.
Figure 6: The ablation result of repair mechanism.
3.5 Computational overhead analysis
Figure 7 compares the token costs between individual- and population-level LLM-based EVO in Netscience. Initial cost refers to the tokens required for a single operation, while total cost includes the additional tokens for repairs. The results show that population-level EVO incurs much lower costs than the individual-level EVO for both crossover and mutation in the entire process, no matter what LLM is used. It can be seen that the repair costs for the population-level approach are higher than those for the individual-based method. Nevertheless, both theoretically and practically, the total costs of the population-level method are lower. As LLMs continue to advance, the accuracy of LLM-based EVO will improve, leading to reduced repair costs, which further show the superiority of population-level optimization.
Figure 7: The computational cost of different cases.
4 Conclusion
To facilitate evolutionary optimization by LLMs, we have conducted a comprehensive evaluation of LLM-based EVO by establishing a set of criteria. Our assessment offers insights and analysis on the strengths and limitations of using LLMs as black-box EVO. Moreover, we have developed an original and novel population-level optimization approach using LLMs, and our proposed repair mechanism to mitigate uncertainties in LLM-generated outputs is also significant. This innovation enhances reliability and has the potential to be extended to numerous practical applications.