### v2

# Can Large Language Models Be Trusted as Evolutionary Optimizers for Network-Structured Combinatorial Problems?

Abstract
Large Language Models (LLMs) have shown impressive capabilities in language understanding and reasoning across diverse domains. Recently, there has been increasing interests in utilizing LLMs not merely as assistants in optimization tasks, but as active optimizers, particularly for network-structured combinatorial problems. However, before LLMs can be reliably deployed in this role, a fundamental question must be addressed: Can LLMs iteratively manipulate solutions that consistently adhere to problem constraints? In this work, we propose a systematic framework to evaluate the capacity of LLMs to engage with problem structures. Rather than treating the model as a black-box generator, we adopt the commonly used evolutionary operators as optimizer and propose a comprehensive evaluation framework that rigorously assesses the output fidelity of LLM-generated operators across different stages of the evolutionary process. To enhance robustness, we introduce a hybrid error-correction mechanism that mitigates uncertainty in LLM outputs. Moreover, we develop a cost-efficient population-level optimization strategy that significantly improves efficiency compared to traditional individual-level approaches. Extensive experiments on a representative node-level combinatorial network optimization task demonstrate the effectiveness, adaptability, and inherent limitations of LLM-based operators. Our findings offer new perspectives on the integration of LLMs into evolutionary computation, providing practical insights for scalable optimization in networked systems.
Index Terms:
Complex networks, combinatorial problems, large language models, evolutionary optimization.
I Introduction
Large Language Models (LLMs), such as GPT (Generative Pre-trained Transformer) trained on vast datasets [1, 2, 3] have the capability to understand and generate human-like text based on the learned patterns. LLMs excel in tasks ranging from simple text completion to complex question answering, demonstrating a nuanced understanding of language context and semantics [4, 5]. The advancement of LLMs [6, 7], has made it possible to revolutionize traditional programming methodologies [8, 9] and reduce human labor. Traditional optimization is labor-intensive and demands domain expertise for precise execution [10]. This challenge intensifies with constrained problems or when tailoring methods for specific needs. Such difficulty is especially pronounced for non-technical users, presenting a major barrier to accessibility and effective application.
There is a growing interest in leveraging LLM not merely as auxiliary tools within optimization pipelines, but as primary solvers [11, 12]. This vision has already been materialized in [13], where optimization of combinatorial problems is achieved not through mathematical analysis or traditional programming but via prompt engineering only. In this work, some combinatorial problems have been used as illustrative examples, such as the traveling salesman problems (TSP) and the solution-score pair is fed into LLMs for the refined solution. Evolutionary optimization has long been recognized as a powerful approach for solving complex problems, valued for both its simplicity and effectiveness. Motivated by these strengths, recent efforts have begun to explore the use of LLMs within evolutionary optimization frameworks. For example, Liu et al. [14] proposed a framework called LMEA, making LLMs function as crossover and mutation operators to solve TSP with up to 20 nodes. Brahmachary et al. [15] leveraged LLMs to optimize the continuous problem with elitism mechanism and demonstrate the potential of LLM-based optimizer. Furthermore, liu et al. [16] utilized LLMs to serve as black-box optimizers for decomposition-based MOEA in a zero-shot manner. In [17], Meyerson et al. explored different structures that LLMs can operate on as crossover, including binary strings, mathematical expressions, text and code.
Despite these advancements, LLMs have shown variable performance across different domains; they excel in some areas such as strategy formulation [18, 19, 20] but not in tasks requiring precise arithmetic and logical reasoning [21, 22]. Therefore, for LLMs to reliably and robustly fulfill this role, a foundational challenge must be addressed: Can LLMs iteratively refine solutions while ensuring that all domain-specific constraints are preserved throughout the optimization process? In combinatorial settings, particularly those involving graphs, schedules, or other structured domains, maintaining feasibility is a challenging and essential requirement. Although encoding schemes differ considerably depending on the task, such as using permutation-based encodings for scheduling problems and binary or graph-based representations for network-related tasks, the fundamental process of manipulating these encodings remains consistent within the evolutionary algorithm framework. Core operations like crossover and mutation are applied universally, typically involving the exchange or replacement of elements, regardless of the specific encoding. Therefore, despite variations in representational strategies, evolutionary algorithms adopt a unified methodology for manipulate the solution space.
To date, there is still a lack of comprehensive investigation to evaluate the effectiveness and reliability of LLMs as optimizers, and to identify the factors affecting their performance. Huang et al. [23] conducted an investigation on LLMs as normal optimizers on different problems, but it is only targeted at one-shot optimization, and some important aspects such as reliability, scalability, and computational cost remain underexplored. Due to the versatility of evolutionary optimizer on the network-related problem, it will serve the illustrative example in our work. Thus, we here aim to conduct a thorough assessment of the performance of LLMs as operators in all stages of evolutionary optimization, some of which, like LLM-based initialization and selection, still remain underexplored in the existing literature. We will also provide insights into the suitability of LLMs, clarifying their capabilities and limitations. Our evaluations compare LLM-based approaches with software-based evolutionary computation to assess the accuracy of LLMs in meeting expected outcomes and their effectiveness in fulfilling customized requirements.
As evolution optimization is an iterative process, an error in any step will yield a cascading effect, leading to a failure of optimization. Recognizing the inherently probabilistic nature of LLMs outputs [24, 25, 26] and the high requirements on the quality of the generated solutions during evolutionary optimization, we develop different stringent sets of standards for LLMs outputs in various levels to rigorously measure solutions in terms of their format, diversity, and conformity to problem constraints. We also introduce a set of corresponding error repair mechanisms with precisely tailored prompts to enhance the reliability of the LLM-based EVO.
In addition, we explore a cost-effective method that utilizes the entire population as optimization unit for the LLM-based reproduction operator and compare this method with another common approach that implements optimization on individual-by-individual. We specifically analyze the quality of generated solutions by these two distinct ways as well as their computational overhead. Some of our findings in this work are summarized as follows:
$\bullet$ LLMs are capable of performing selection, crossover, and mutation operations within evolutionary optimization. However, their effectiveness is highly sensitive to hyperparameters such as population size and solution length, which must be carefully tuned to maintain performance.
$\bullet$ LLMs can effectively perform decision-making tasks such as selection, often outperforming traditional heuristic methods in adaptability. Unlike fixed-rule heuristics, LLMs can incorporate complex contextual information such as fitness values, element rankings when making selection decisions.
$\bullet$ The effectiveness and reliability of LLM-based evolutionary optimization are closely tied to the capacity of the underlying language model, particularly its number of parameters. Larger models generally demonstrate stronger reasoning abilities, greater contextual understanding, and improved consistency across generations.
$\bullet$ The initialization phase can be computationally intensive and less suitable for LLMs. As dataset size increases, LLM performance in this phase often degrades significantly, suggesting a need for auxiliary strategies or preprocessing.
$\bullet$ LLMs are sensitive to the volume and complexity of input data. When exposed to large-scale inputs, they are prone to generating infeasible or suboptimal solutions, highlighting the importance of integrated correction and repair mechanisms throughout the optimization process.
$\bullet$ Population-level LLM-based EVO offers greater computational efficiency compared to individual-level approaches. By operating on the entire population in a single prompt, it reduces the number of model calls and repetitive descriptions of operations.
$\bullet$ LLM-based reproduction is highly scalable in addressing network-structured problems with respect to token cost. This efficiency stems from the use of compact element ID representations, which scale logarithmically with dataset size, allowing LLMs to handle large problems more gracefully than if inputs scaled linearly.
II Related work
In this section, we will review the existing literature on the Synergy of LLM and optimization, and the application of evolutionary optimization on network-structured problems.
II-A Synergy of LLMs and optimization
Yu and Liu [27] comprehensively investigated the synergy of LLMs and evolutionary optimization and discussed different-angle applications. Wu et al. classified existing works regarding the synergy of LLMs and evolutionary computation into two main branches [10]: LLM-based black-box search operators [28] and LLM-based optimization algorithm automation [29, 30].
LLMs can potentially automate and optimize the design of evolutionary operators as indicated in [12], thereby reducing the need for manual tuning and domain-specific adjustments such as [31]. Yang et al. proposed to use LLMs as optimizer named Optimization by PROmpting (OPRO) to solve the traveling salesman problem [13], in which the previously generated solution and its evaluation value are used as part of the prompt for next generation. In the study by Meyerson [17], LLMs are employed as crossover operators to derive new solutions from parental inputs. Brownlee [32] also presented LLMs effectively functioning as mutation operators that enhance the search process. Liu et al. introduced a novel framework known as LLM-driven EA (LMEA) [14], which utilizes LLMs for both crossover and mutation operations. This approach highlights the adaptability of LLMs, where search behaviors can be easily modified by adjusting the LLM temperatures. Furthermore, LLM-based search operators can also be adapted for multi-objective scenarios by segmenting traditional optimization tasks into sub-problems [16]. In [33], Wang et al. explored the applicability of LLMs on constrained multiobjective optimization problems and achieved the promising results compared to the existing methods.
LLMs have shown significant potential in automated evolutionary optimization, autonomously optimizing, creating, and improving algorithms to effectively tackle optimization challenges [34]. As demonstrated in [19], heuristic searched by LLMs achieves excellent performance in different complex combinatorial problems. In [35], Liu et al. proposed a new method called Algorithm Evolution using the Large Language Model (AEL) that directly considers algorithms as individuals for evolution. AEL was then further extended to the design of guided local search algorithms [36], showing the strength of the LLM-based method over human-designed algorithms. After that, they extended AEL to an advanced model called Evolution of Heuristics (EoH) by exploring various prompts to solve different combinatorial tasks. LLMs analyze swarm intelligence algorithms to obtain a hybrid algorithm that combines various strengths of existing methods [37]. Mao et al. explored LLM-enhanced algorithm automation for identifying critical nodes [38]. In this approach, various heuristics are initialized as populations and then evolve with the assistance of LLMs. Morris et al. [29] introduced a thought reflection method to guide the LLMs-based algorithm automation and improved the accuracy of ExquisiteNetV2 models. The evolution process was also studied to enhance the adaptability and convergence by [39], where the mutation rate is adaptively adjusted with dynamic prompt.
LLMs can also be used as a surrogate model with the help of in-context learning to efficiently analyze the quality of the solution, as shown in [40]. Furthermore, LLMs can also assist in algorithm selection, as demonstrated in [41], through analyzing the code to grasp both its structural and semantic elements, along with the contextual understanding. An emerging direction worth noting involves using evolutionary optimization to search for the optimal prompt, enabling LLMs to achieve excellent performance. Notable examples of this approach can be found in the work of [42, 43, 44].
II-B Metaheuristic optimization on network-structured problems
Network-structured combinatorial problems play a crucial role in various practical domains due to their complexity and widespread applicability [45, 46]. The utility of evolutionary optimization in addressing discrete and non-linear problems has significantly facilitated its adoption across diverse fields, particularly in the context of complex networks [47, 48] and various combinatorial tasks [49, 50]. These optimization algorithms excel due to their inherent capability to navigate complex solution spaces effectively, making them suitable for tasks such as truck scheduling [51, 52] and job shop scheduling [53, 54]. In complex network, evolutionary optimization addresses several intricate combinatorial challenges, including influence maximization [55, 56], robustness analysis [57, 58], and community deception [59, 60]. Such tasks pose particular difficulties due to their non-linear interactions and discrete search spaces, factors that evolutionary algorithms are particularly adept at handling. For instance, Zhang et al. leveraged multi-objective evolutionary optimization to efficiently identify local communities within complex networks. Similarly, Liu et al. [61] presented an evolutionary framework employing a relationship-driven strategy to identify critical node subsets, achieving significant improvements across multiple benchmarks. Notably, evolutionary algorithms have also been effectively applied in network-related tasks such as important node identification [62], community discovery [63, 64], network reconstruction [65, 66], and network module recognition [67, 68]. These applications underscore the versatility and robustness of evolutionary optimization methodologies in addressing complex network problems.
A critical aspect of evolutionary optimization frameworks is the encoding scheme, which translates a problem into a suitable representation for algorithmic manipulation. While encoding schemes vary significantly based on the nature of the tasks, such as permutation-based encodings for scheduling tasks and binary or graph-based encodings for network tasks, the underlying mechanism of manipulating these representations remains consistent within the evolutionary algorithmic framework. Operations such as crossover and mutation are universally applied, regardless of the encoding specifics in the same way, i.e., exchange and replace elements. Thus, despite differences in representational strategies, evolutionary algorithms maintain a unified approach to exploring and exploiting the solution space.
III LLM-based evolutionary optimizer
In the combinatorial problems in complex networks, the solution representation within the evolutionary optimization framework is generally defined as
$[\text{X}_{1},\text{X}_{2},...,\text{X}_{n}],$ (1)
where X depends on the specific task and $n$ is the solution size [69]. In this work, we choose the influence maximization [70, 71] as an illustrative example. The problem is defined as: Given a graph $G=(V,E)$, where $V$ represents the set of nodes and $E$ represents the set of edges, the objective is to find a subset of nodes $S\subseteq V$ that maximizes the influence across the network. As for network-related problems, the solution representation can simply be
$[\text{Node}_{1},\text{Node}_{2},...,\text{Node}_{n}].$ (2)
Note that our focus is not on solving any specific network-related problem but on evaluating the performance of LLMs as operators for combinatorial problems in complex networks. The reason for selecting this problem is that it is general enough to provide insights and observations for LLMs as EVO due to the representation similarity on combinatorial problems. To enhance generality, ‘element’ will replace ‘node’, and ‘dataset’ will replace ‘graph’ in the rest of the discussion. For the selection, the input to LLMs is
$\{([\text{X}_{1}^{(1)},\ldots,\text{X}_{n}^{(1)}],f_{1}),\ldots,([\text{X}_{1}^{(k)},\ldots,\text{X}_{n}^{(k)}],f_{k})\},$ (3)
where $k$ is the population size and $\text{X}_{i}^{(j)}$ refers to the $i$-th element in the $j$-th solution in the population. It can be simplified as
$\{(S_{1},f_{1}),...,(S_{k},f_{k})\},$ (4)
where $S_{i}$ is the index of solution $[\text{X}_{i}^{(1)},\ldots,\text{X}_{n}^{(i)}]$ and $f_{i}$ is its fitness value.
Let $\mathbf{F}$ be combination of LLM-based evolutionary operators, i.e.,
$\mathbf{F}=\{\mathbf{F}_{I},\mathbf{F}_{S},\mathbf{F}_{C},\mathbf{F}_{M}\},$ (5)
where $\mathbf{F}_{I}$, $\mathbf{F}_{S}$, $\mathbf{F}_{C}$ and $\mathbf{F}_{M}$ indicate LLM-based initialization, selection, crossover, and mutation, respectively.
Figure 1: (A) The diagram of LLM-based EVO with the proposed validation and repair mechanism. All four phases in evolution optimization along with repair are based on LLMs. (B) The illustration of population- and individual-level LLM-based EVO. (C) An example of errors encountered in population-level LLM-based crossover. The error message (if any), customized corrected prompt, and previous deficient output will be provided to LLMs for repair.
The first phase, Initialization is defined as $\mathcal{P}_{0}=\mathbf{F}_{I}(G)$, where $\mathcal{P}_{0}$ refers to the initialized population and $G$ refers to the input data. A strategic reduction in search space can speed up convergence and improve effectiveness [72]. However, based on current research [21], it is not feasible to feed datasets directly to LLMs, either in the form of adjacency matrices or text descriptions, because the reasoning ability of current LLMs on graphs remains limited. Therefore, instead of providing LLMs with just a dataset, we opt to input element IDs along with easily accessible metrics. Following this, LLMs will be instructed to (1) Rank the elements based on a specific metric; (2) Select a percentage of top elements as candidates; and (3) Sample filtered elements to form the initialized population. Several basic restrictions must be followed to ensure the quality of the initialized population: (1) Each solution in the population should contain the same number of elements with no duplicates; (2) The solutions within the population should be as diverse as possible; and (3) elements outside the candidate set are not allowed to appear in the population.
The Selection phase in evolutionary optimization is a crucial step where individuals from the current population are selected as the next generation. This is represented as $\mathcal{P}_{l+1}=\mathbf{F}_{S}(\mathcal{P}_{l},f(\mathcal{P}_{l}))$, where $\mathcal{P}_{l+1}$ refers to the population after applying operators of $\{\mathbf{F}_{C},\mathbf{F}_{M}\}$ on $\mathcal{P}_{l}$ and $f(\mathcal{P}_{l})$ represents the fitness values of solutions in $\mathcal{P}_{l}$. Extensive studies on selection strategies, such as roulette wheel selection, aim to preferentially choose individuals with higher fitness scores. This operation allows the population to evolve towards an optimal solution over successive generations. In $\mathbf{F}_{S}$, we will not rely on any specified strategy but instead follow the common principles: (1) Solutions with low fitness are not allowed to be selected; and (2) Each solution can be selected multiple times but not excessively, to maintain diversity.
As for reproduction, the individual-level LLM-based operators of Crossover $\mathbf{F}_{C}^{S}$ and Mutation $\mathbf{F}_{M}^{S}$ are
$\begin{split}&\mathcal{P}_{l+1}=\bigcup_{\mathcal{S}_{i},\mathcal{S}_{j}\in\mathcal{P}_{l}}\mathbf{F}_{C}^{S}(\mathcal{S}_{i},\mathcal{S}_{j}),\\ &\mathcal{P}_{l+1}=\bigcup_{\mathcal{S}\in\mathcal{P}_{l}}\mathbf{F}_{M}^{S}(\mathcal{S}),\end{split}$ (6)
where $\mathcal{S}$ denotes the solution in population $\mathcal{P}$. The individual-level operator is the conventional method for implementing evolutionary optimization [14, 16]. While this individual-level method provides a precise and controlled way to reproduce the population, it may suffer from two problems: (i) Handling solutions individually or in pairs might not scale efficiently as the population size grows in the framework of LLM-based EVO; and (ii) LLMs may lack broader contextual information about the population’s overall diversity, which could impede their ability to optimize effectively. To resolve these critical issues, we propose here a new population-level optimization, defined as follows:
$\begin{split}&\mathcal{P}_{l+1}=\mathbf{F}_{C}^{P}(\mathcal{P}_{l}),\\ &\mathcal{P}_{l+1}=\mathbf{F}_{M}^{P}(\mathcal{P}_{l}).\\ \end{split}$ (7)
Here we distinguish between these two methods: All evolutionary optimizations rely on populations, typically selecting one or a pair of individuals from the population iteratively to input into LLMs, which is referred to as individual-level in our work. In contrast, our proposed population-level method is to input the entire population into LLMs for optimization.
One unique advantage of the proposed population-level LLM-based optimization lies in its global awareness of the diversity of the entire population. This enables the potential for intelligent reproduction, specifically selecting individuals most likely to produce optimal solutions for recombination, rather than relying on random selection or investing substantial effort in designing complex algorithms to choose individuals, as is typical in individual-level reproduction, whether based on code or LLMs.
The four components, initialization, selection, crossover, and mutation are all achieved by LLMs. The pseudocode of population- and individual-level LLM-based EVO is shown in Algorithms 1 and 2 respectively. The key difference is in line 7 of Algorithm 2, where solutions from the population are iteratively selected for reproduction, unlike in population-level optimization, which does not require this step. This generality is also reflected in our prompt engineering where little network domain knowledge is needed (please refer to Tables III and IV for LLM-based mutation).
Algorithm 1 Population-level LLM-based EVO with repair mechanism
Input: Dataset $G$, fitness function $f$, maximum number of generations $N_{\text{max}}$
Output: Optimized node set for influence maximization
1:Initialize population $\mathcal{P}=\{\mathcal{S}_{1},\mathcal{S}_{2},\cdots,\mathcal{S}_{\mathcal{N}_{P}}\}$ instructed by LLMs
2:Check and (repair) the output
3:while iteration count $<N_{\text{max}}$ do
4:     Calculate fitness value of solutions in $\mathcal{P}$: $f(\mathcal{P})$
5:     Select solutions as $\mathcal{P}$ for reproduction instructed by LLMs
6:     Check and (repair) the output
7:     Perform crossover on population instructed by LLMs
8:     Check and (repair) the output
9:     Perform mutation on $\mathcal{P}$ instructed by LLMs
10:     Check and (repair) the output
11:end while
12:Return solution with the highest fitness value in the population.
Algorithm 2 Individual-level LLM-based EVO with repair mechanism
Input: Dataset $G$, fitness function $f$, maximum number of generations $N_{\text{max}}$
Output: Optimized node set for influence maximization
1:Initialize population $\mathcal{P}=\{\mathcal{S}_{1},\mathcal{S}_{2},\cdots,\mathcal{S}_{n}\}$ instructed by LLMs
2:Check and (repair) the output
3:while iteration count $<N_{\text{max}}$ do
4:     Calculate fitness value of solutions in $\mathcal{P}$: $f(\mathcal{P})$
5:     Select solutions as the population for reproduction instructed by LLMs
6:     Check and (repair) the output
7:     for $S\in\mathcal{P}$ do
8:         Perform crossover on $S$ instructed by LLMs
9:         Check and (repair) the output
10:         Perform mutation on $S$ instructed by LLMs
11:         Check and (repair) the output
12:     end for
13:end while
14:Return solution with highest fitness value in the population.
III-A Computational cost of population- and individual-level LLM-based EVO
As we aim to assess the fidelity of LLM-based EVO, the benchmark is based on precise coding methods. It is not applicable to compare their computational costs and efficiency since LLMs rely on online resources (e.g., token count) while code-based methods depend on local hardware. Then, we will discuss the difference between population- and individual-level optimization regarding computational cost.
Crossover: Let $\mathcal{V}$ be the token cost of LLMs, and $\mathcal{V}(P)$ and $\mathcal{V}(S)$ refer to the token of the entire population and individual solution, respectively. The input to LLMs consists of prompt $\mathcal{T}_{C}^{S}$ ($\mathcal{T}_{C}^{P}$) and a pair of solutions (population), where $\mathcal{V}(\mathcal{T}_{C}^{S})\approx\mathcal{V}(\mathcal{T}_{C}^{P})$. Note that $\mathcal{V}(\mathcal{T}_{C}^{P})$ is slightly higher than $\mathcal{V}(\mathcal{T}_{C}^{S})$ because there will be an additional text like ”Please randomly select pairs of solutions and continue applying the crossover operation until the number of newly created solutions matches the population size.” in the population-level method, in which the token cost is $\mathcal{V}(\mathcal{T}_{C}^{P})+\mathcal{V}(P)$. In contrast, the individual-level method requires LLMs to implement crossover operation on pair by pair, which costs $(\mathcal{V}(\mathcal{T}_{C}^{S})+2*\mathcal{V}(S))*\frac{\mathcal{N}_{p}}{2}$ where $\mathcal{N}_{p}$ refers to the population size. As $\mathcal{V}(P)=\mathcal{V}(S)*\mathcal{N}_{p}$, the cost difference between individual- and population-level crossover is $\mathcal{V}(\mathcal{T}_{C}^{S}*\mathcal{N}_{p}/2)-\mathcal{V}(\mathcal{T}_{C}^{P})\approx\frac{\mathcal{N}_{p}-2}{2}*\mathcal{V}(\mathcal{T}_{C}^{S})$.
Mutation: As for individual-level mutation, it will run $\mathbf{F}_{M}^{S}$ for $\mathcal{N}_{p}$ times since it will mutate the solution individual by individual. Thus, its cost difference with population-level mutation is $\mathcal{V}(\mathcal{T}_{M}^{S}*\mathcal{N}_{p})-\mathcal{V}(\mathcal{T}_{M}^{P})\approx(\mathcal{N}_{p}-1)*\mathcal{V}(\mathcal{T}_{M}^{S})$. Clearly, the proposed population-level method can save roughly $\frac{\mathcal{N}_{p}-2}{2}*\mathcal{V}(\mathcal{T}_{C}^{S})+(\mathcal{N}_{p}-1)*\mathcal{V}(\mathcal{T}_{M}^{S})$ in each round as compared to individual-level method, and this advantage in cost will be further amplified when population size and evolution round are set to a high number.
Error Type Error Index Error Description
Format Error E1 The output is not in the required format.
E2 The output contains non-integer elements.
Critical Error E3 The selected candidates significantly deviate from the ground truth.
E4 The size of candidates falls significantly short of meeting the requirements.
E5 The size of the population falls significantly short of meeting the requirements.
E6 The selected population contains one solution too many times.
E7 Any solution appears in the population where all elements are the same.
E8 The number of different elements in the solution changes significantly.
E9 The number of different solutions in the population changes significantly.
Moderate Error E10 The size of some solutions fails to meet the requirement.
E11 The solution (population) after the operation remains the same as before.
E12 The size of the population fails to meet the requirement.
E13 The new solution contains duplicated elements.
E14 The new solution contains invalid elements not found in candidate nodes.
E15 The selected population contains those with very low fitness.
TABLE I: The errors of LLM outputs during evolutionary optimization.
III-B Error repair
Phase Format Error Critical Error Moderate Error
E1 E2 E3 E4 E5 E6 E7 E8 E9 E10 E11 E12 E13 E14 E15
Candidate selection - - - - - - - - - - -
Initialization - - - - - - - -
Selection - - - - - - - - - -
Crossover (P) - - - - -
Crossover (S) - - - - - - - -
Mutation (P) - - - -
Mutation (S) - - - - - - -
TABLE II: The required validations for the LLM output in different phases of evolutionary optimization.
The probabilistic nature of LLMs can lead to occasional undesirable results. As evolutionary optimization is an iterative process, an error early in the process can ‘blow up’ in subsequent cycles, creating a cascading effect to ruin the optimization process. To investigate the reliability of LLM-generated outputs across different phases, we establish rigorous validation standards and systematically investigate potential errors encountered during simulation. These errors are categorized into format errors and quality errors, with the latter further divided into critical and moderate types. Specifically,
  • Format Error refers to generated solutions that fail to adhere to required structures.
  • Critical Error refers to solutions that, while structurally correct, may cause system failures, demanding immediate resolution to prevent extensive damages.
  • Moderate Error refers to solutions that can cause minor disruptions but generally will not lead to complete failure of optimization.
The summary of the different types of errors are listed in Table I and details are introduced as follows.
III-B1 Format Error
The format error refers to those that can directly interrupt the running of a program. Since format errors are rare and difficult to fix, it is usually more efficient to request a new generation.
  • E1: The output is not in the required format.
    • Example: When we ask LLMs to generate a list in Python, we expect only that list without any explanation and text for further operations. However, LLMs sometimes add extra text, like ”You are doing a crossover, and the result is …”.
    • Impact: This error can disrupt the optimization process because algorithms depend on a specific data structure.
  • E2: The output contains non-integer elements.
    • Example: The combinatorial problems require discrete integer values while LLMs sometimes output a solution containing non-integer numbers such as [3, 5.9, 2.1, 8].
    • Impact: Non-integer outputs can lead to invalid candidate solutions that cannot be evaluated, thus this error will also interrupt the process as E1.
III-B2 Critical Error
Most critical errors undermine diversity, which is essential for successful optimization. Note that once diversity is severely damaged, then it is very difficult to repair as only mutation contributes to exploration in the entire evolutionary optimization. Therefore, this kind of error is not allowed in the optimization. In a manner consistent with addressing format errors, a new output will be requested instead of attempting to correct an unacceptable one.
  • E3: The selected candidates significantly deviate from the ground truth.
    • Example: We input all possible elements, each with a specific metric, and seek the top 50% (for example). However, LLMs outputs may only minimally overlap with the ground truth, indicating that most of their suggestions are not suitable candidates.
    • Impact: This case will reduce the quality of the initialized population, further impacting subsequent optimization.
  • E4: The size of candidates falls significantly short of meeting the requirements.
    • Example: We input all possible elements, each with a specific metric, and seek the top 50%. However, LLMs may output only 10% elements.
    • Impact: Such a number is insufficient to guarantee the diversity of the population as the solutions in the initialized population will be very similar.
  • E5: The size of the population falls significantly short of meeting the requirements.
    • Example: We input the entire population consisting of 30 solutions into LLMs but only receive 10 solutions as output.
    • Impact: The drastic reduction in population size will severely affect the diversity of the population. The population size may be restored to the predefined number but the diversity remains poor as there will be a lot of repetitive solutions.
  • E6: The selected population contains one solution too many times.
    • Example: We input a set of solution IDs and wish to filter those low-fitness solutions. However, LLMs may occasionally produce repetitive IDs for numerous identical high-fitness solutions.
    • Impact: Although the low-fitness solutions are filtered, over-selecting a single solution damages the diversity of the population.
  • E7: Any solution appears in the population where all elements are the same.
    • Example: LLMs sometimes return several solutions like $[A_{1},A_{1},A_{1},A_{1}]$ but the constraint is that no repetitive element is allowed within the same solution.
    • Impact: When this kind of error occurs, we discovered that the number of erroneous solutions in the population is not a few but a lot. After several rounds of crossover, all solutions of the entire population will be polluted, resulting in poor diversity.
  • E8: The number of different elements in the population changes significantly.
    • Example: When we input a population containing 50 different elements, LLMs sometimes will output a population containing only 20 elements.
    • Impact: The drastic reduction in the number of different elements will severely affect the diversity of the population.
  • E9: The number of different solutions in the population changes significantly.
    • Example: This error is similar to E6 that occurs during selection, whereas this error occurs during reproduction. When we input a population consisting of 30 various solutions to LLMs, the output LLMs may also contain 30 solutions but most of which may be identical.
    • Impact: The drastic reduction in the number of different solutions will severely affect the diversity of the population.
III-B3 Moderate Error
The moderate errors have a high chance of being repaired during the optimization thus they will not yield severe impact.
  • E10: The size of some solutions fails to meet the requirement.
    • Example: When we input a solution consisting of 10 elements to LLMs for reproduction, the output may have 9 or 11 elements.
  • E11: The solution (population) after the operation remains the same as before.
    • Example: When we input a solution (population) to LLMs for reproduction, the output remains unchanged. For example, the solution $[A_{1},A_{2},A_{3},A_{4}]$ is still $[A_{1},A_{2},A_{3},A_{4}]$ after mutation.
  • E12: The size of the population fails to meet the requirement.
    • Example: When we input a population consisting of 30 solutions for reproduction, LLMs may output a population containing 28, 29, 31, or 32 solutions.
  • E13: The new solution contains duplicated elements.
    • Example: The issue resembles E7, but is less severe and involves only 2 or 3 identical elements in the output solution.
  • E14: The new solution contains invalid elements not found in candidate nodes.
    • Example: Suppose we have 100 elements and filter the 50 as candidates, the LLMs sometimes output a solution containing the element in the other 50. We found that LLMs did not produce elements outside these 100, thus we categorize this error as moderate.
  • E15: The selected population contains those with very low fitness.
    • Example: This error is like E11 but this is for selection. When we input a list of solutions to LLMs for selection, their output will be the same as what we input before.
An example of possible errors encountered in the population-level crossover is given in Figure 1(C). The outputs generated in different phases will undergo different validations due to the varying requirements for each phase. In addition, the output requirements are different when the input is an individual solution (denoted as $S$) and the entire population (denoted as $P$). The detail is shown in Table II.
During optimization, particularly in environments with numerous constraints, managing invalid solutions is critical in ensuring successful optimization. To this end, we introduce a robust repair mechanism that maintains solution feasibility and ensures that each iteration contributes positively toward finding an optimal solution. Let $\mathcal{O}$ be the output from $\{\mathbf{F}_{I},\mathbf{F}_{S},\mathbf{F}_{C},\mathbf{F}_{M}\}$, the refined output is defined as
$\mathcal{O}\leftarrow\mathbf{F}_{R}(\mathcal{O},E_{x},R_{x}),$ (8)
where $\mathbf{F}_{R}$ refers to the LLM-based repair operator. $E_{x}$ and $R_{x}$ denote the error message and targeted repair prompt.
During the optimization, each output undergoes the three aforementioned examinations. If any error is detected, the repair mechanism is triggered. We will check and repair each type of error individually. Format errors and critical errors are particularly detrimental. The former interrupts the entire optimization process, and the latter can heavily corrupt the population, diverting the optimization from its optimal path. We will avoid repairing these two types of errors due to the complexity involved. Instead, a new generation is directly requested to obtain a valid solution. Due to their destructive impact, solutions failing these checks cannot be used and the previous phase’s solution will be used for the next phase.
In contrast, solutions with moderate errors that are not successfully repaired are allowed to enter into the next phase. This approach is adopted because (1) the error can always be mitigated after repair; and (2) moderate errors do not yield cascading effects on the entire optimization, and there will still be an opportunity to repair them in the next phases. For the process of validation and repair, please refer to Algorithm 3. In each stage, the output of LLMs will undergo the corresponding check and repair (if any). For the checklist of each stage, please refer to Table 2 of the main text.
Algorithm 3 CheckAndRepair
1:Output $\mathcal{O}_{x}$ obtained by $\mathbf{F}_{x}\in\{\mathbf{F}_{I},\mathbf{F}_{S},\mathbf{F}_{C},\mathbf{F}_{M}\}$
2:Refined output
3:Retrieve the checklist $\mathbf{L}_{F}$ in the phase regarding $\mathbf{F}_{x}$
4:for each $E_{x}$ in $\mathbf{L}_{F}$ do
5:     Check the input regarding $E_{x}$
6:     if $E_{x}$ passed then
7:         $\mathcal{O}_{x}\leftarrow\mathcal{O}_{x}$triangleright Output $\mathcal{O}_{x}$ remains unchanged
8:     else
9:         Retrieve the corresponding repair prompt $R_{x}$
10:         $\mathcal{O}_{x}\leftarrow\mathcal{F}_{R}(\mathcal{O}_{x},E_{x},R_{x})$triangleright pass the repaired result to the next check
11:     end if
12:end for
13:Return the final output $\mathcal{O}_{x}$
[…]
TABLE III: The prompt of population-level mutation
[…]
TABLE IV: The prompt of individual-level mutation
III-C Solution evaluation
After the output undergoes the check and, if necessary, repair process, we categorize it into the following cases:
  • Approved ($\mathsf{Q}_{\mathsf{app}}$): The solution meets all requirements and standards perfectly, with no errors or deficiencies. It is ready for implementation without any modifications.
  • Repaired ($\mathsf{Q}_{\mathsf{rep}}$): The solution had minor issues that did not meet the necessary standards, but these have been addressed, and it now meets the required criteria.
  • Acceptable ($\mathsf{Q}_{\mathsf{acc}}$): The solution contains flaws that fail to be completely corrected, but it still functions adequately and meets the minimum necessary criteria for use, albeit not optimally.
  • Rejected ($\mathsf{Q}_{\mathsf{rej}}$): The solution has format or critical error even though after an attempt of repair, rendering it completely unusable.
For the format and critical errors, the acceptable case $\mathsf{Q}_{\mathsf{acc}}$ is not applicable due to their destructive impact on the population. If a solution encounters these errors and cannot be repaired, it must be rejected. In this case, the output with moderate error from last check will pass through to the next phase if repair fails.
IV Experimental studies
In this section, we will examine the ability of LLMs to manipulate the solution of the network-structured problems.
IV-A Experimental settings and dataset
The fidelity and reliability of LLM-based EVO of different stages are validated on various datasets in different settings. The population size and the evolution round are set to 30. The number of tests for candidate selection is 100, and the results are averaged from 10 independent simulations.
We investigated using the population as the fundamental optimization unit, where all solutions undergo crossover and mutation. To fairly compare the population and individual-level optimization by LLM, we thus set the crossover and mutation rate to 1.0 for individual-level optimization. The ablation study employs GPT-4o, with initialization and selection done using GPT-4, while other tests utilize both GPT-4 and GPT-3.5. Other experimental settings are shown in Table V.
Parameter Value
LLM model GPT-3.5/4/4o
LLM temperature 0.8
Maximum token 1400
Population size 30
Number of generations per simulation 30
Solution size 10
Crossover rate 1.0
Mutation rate 1.0
Candidate number for initialization 50%
Node metric for candidate selection Degree
TABLE V: The experimental setting of LLM-based EVO.
Table VI provides more details about the tested networks, including $|V|$ and $|E|$ for the number of nodes and edges, respectively, $\langle K\rangle$ for the average degree, and CC and ASD for the clustering coefficient and average shortest distance. The data is available online [^1] http://www-personal.umich.edu/mejn/netdata..
Network $|V|$ $|E|$
Dolphins 62 159
Netscience 379 914
Erods 433 1,314
Email 1,005 25,571
Astro 14,845 119,652
TABLE VI: Topological information of networks.
IV-B Optimization problem formulation
The broad utility of evolutionary optimization for tackling discrete, non-linear problems has led to their extensive adoption in combinatorial challenges [73, 74]. In this study, we use the problem influence maximization that is extensively explored by the evolutionary computation community as an illustration [75, 57, 76] due to its generalizability. Given a graph $G=(V,E)$, where $V$ represents the set of nodes and $E$ represents the set of edges, the objective is to find a subset of nodes $S\subseteq V$ that maximizes the influence across the network. Let $\{C_{1},C_{2},\ldots,C_{k}\}$ be the communities partition. For each community $C_{i}$, we can evaluate the influence of seed node set $S$ by calculating the intersection of their influenced nodes $I(S)$ with the nodes in $C_{i}$ as follows:
$I_{i}(S)=|I(S)\cap C_{i}|,$ (9)
where $I(S)$ refers to the function that simulates the spread of influence from the initial set $\mathcal{S}$.
The weight of each community $w_{i}$ is proportional to its size relative to the network. The fitness function $F(S)$ is then calculated as the weighted sum of the influences within each community, i.e.,
$f(S)=\sum_{i=1}^{k}\frac{|C_{i}|}{|V|}\cdot|I(S)\cap C_{i}|.$ (10)
IV-C Optimization validation of LLM-based EVO
IV-C1 Initialization
Table VII shows the results of the pass ratio of format and critical check in the candidate selection, consisting of ranking and filtering. As seen, LLMs effective on small datasets fail to generate valid outputs for large datasets. For example, on Dolphins, LLMs achieve 100.0% $\mathsf{Q}_{\mathsf{acc}}$ on both checks. We compare it with larger networks and observe an obvious decrease in pass ratios; particularly, the critical check pass drops to about 40%. Thus, it can be seen that LLMs’ performance in calculation tasks, like candidate selection, will decline as the amount of input data increases.
While LLMs possess the potential for global awareness, enabling them to generate a more diverse and well-distributed initial population, practical limitations arise when applied to large-scale networks. To make informed decisions, LLMs require structured input that includes node IDs along with relevant metrics, such as degree, betweenness centrality, or clustering coefficients. However, encoding this information for all nodes in a large graph can quickly exceed the model’s token limit, making full-graph initialization infeasible. This token overhead significantly restricts scalability, necessitating pre-filtering strategies or metric-based candidate selection before LLM engagement.
Dataset Dolphins Netscience Erods
Format check pass 100.0% 95.0% 94.0%
Critical check pass 100.0% 41.0% 42.0%
TABLE VII: The validation of LLM-based candidate selection.
Therefore, we only test the LLM-based initialization in the small networks as a preliminary study. The LLM-based initialization shows superior fitness at the starting point and during the optimization compared to random initialization, as shown in Figure 2. This result shows that LLMs can effectively handle population initialization, reducing human effort for networks of hundreds of nodes.
Figure 2: The comparison of different initialization methods.
IV-C2 Selection
The following are selection strategies for the comparison with LLM-based selection. Random selection chooses individuals randomly as parents for the next generation, without considering their fitness. No selection means every individual in the population survives to the next generation, regardless of fitness. Roulette selection assigns selection probabilities based on fitness, with fitter individuals having a higher chance of being selected. Tournament selection involves selecting a subset of individuals randomly and choosing the best from this subset as a parent. The process is repeated as needed.
Figure 3 compares the selection performance of LLM-based selection with other selection strategies. The LLM-based selection demonstrates the highest fitness across the entire optimization generation, with only the tournament strategy achieving comparable performance. It can be deduced that LLMs are effective in decision-making tasks such as in the selection phase.
Figure 3: The comparison of different selection methods.
IV-C3 Reproduction
For fairness, we enhanced the baseline code-based EVO shown in Figure 4 by ensuring it does not violate constraints. As observed, the LLM-based EVO can achieve nearly the same performance as software-based evolutionary optimizers, reflected by their similar fitness. Additionally, the choice of model significantly impacts performance, as evidenced by the poor results of the population-level LLM-based EVO by GPT-3.5.
Figure 4: The comparison of different optimizers.
Figure 5: The validation of moderate error in different phases.
IV-D Reliability analysis
Netscience Format Error (GPT-4) Critical Error (GPT-4) Format Error (GPT-3.5) Critical Error (GPT-3.5)
$\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$
Initialization 100.0% 0.0% 0.0% 97.0% 2.0% 1.0% 98.0% 2.0% 0.0% 84.0% 16.0% 0.0%
Selection 100.0% 0.0% 0.0% 99.3% 0.7% 0.0% 100.0% 0.0% 0.0% 95.0% 4.0% 1.0%
Crossover (P) 99.6% 0.0% 0.4% 97.6% 2.3% 0.0% 95.6% 4.4% 0.0% 78.0% 13.6% 8.4%
Mutation (P) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 98.0% 2.0% 0.0% 92.3% 3.3% 4.4%
Crossover (S) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 99.6% 0.4% 0.0% 100.0% 0.0% 0.0%
Mutation (S) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 96.0% 3.7% 0.3% 99.9% 0.0% 0.1%
TABLE VIII: The validations for the output generated in different phases. The tested network is Netscience. (P) refers to population-level reproduction and (S) refers to individual-level reproduction. Initialization refers to the sampling procedure.
Erods Format Error (GPT-4) Critical Error (GPT-4) Format Error (GPT-3.5) Critical Error (GPT-3.5)
$\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$
Initialization 100.0% 0.0% 0.0% 98.0% 2.0% 0.0% 92.0% 8.0% 0.0% 78.0% 14.0% 8.0%
Selection 100.0% 0.0% 0.0% 92.2% 5.5% 2.3% 100.0% 0.0% 0.0% 91.1% 6.9% 2.0%
Crossover (P) 99.3% 6.3% 0.1% 98.6% 1.4% 0.0% 93.3% 6.3% 0.4% 74.6% 13.4% 12.0%
Crossover (S) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 99.9% 0.1% 0.0% 100.0% 0.0% 1.0%
Mutation (P) 100.0% 0.0% 0.0% 99.3% 0.7% 0.0% 98.0% 1.6% 0.4% 93.0% 3.4% 3.6%
Mutation (S) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 98.0% 1.6% 0.4% 97.0% 1.3% 1.7%
TABLE IX: The validations for the output generated in different phases. The tested network is Erods. (P) refers to population-level reproduction and (S) refers to individual-level reproduction. Initialization refers to the sampling procedure.
Email Format Error (GPT-4) Critical Error (GPT-4) Format Error (GPT-3.5) Critical Error (GPT-3.5)
$\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$
Initialization 100.0% 0.0% 0.0% 96.0% 4.0% 0.0% 98.0% 2.0% 0.0% 94.0% 6.0% 0.0%
Selection 100.0% 0.0% 0.0% 92.2% 5.5% 2.3% 100.0% 0.0% 0.0% 91.1% 6.9% 2.0%
Crossover (P) 98.0% 2.0% 0.0% 94.3% 4.3% 1.4% 96.3% 3.7% 0.0% 71.6% 10.4% 18.0%
Crossover (S) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 99.7% 0.2% 0.1% 100.0% 0.0% 1.0%
Mutation (P) 99.3% 0.6% 0.0% 99.7% 0.3% 0.0% 96.6% 3.4% 0.0% 89.8% 3.6% 6.6%
Mutation (S) 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 94.4% 4.9% 0.7% 100.0% 0.0% 0.0%
TABLE X: The validations for the output generated in different phases. The tested network is Email. (P) refers to population-level reproduction and (S) refers to individual-level reproduction. Initialization refers to the sampling procedure.
Table VIII shows that the ratio of $\mathsf{Q}_{\mathsf{app}}$ regarding format error is consistently above 95% in all phases. Meeting the requirements of critical errors is more challenging as more criteria are involved. GPT-4 performs exceptionally well, achieving a $\mathsf{Q}_{\mathsf{rej}}$ below 1%. Even when issues are occasionally detected, they can be repaired, as seen in the initialization case. In contrast, LLM-based EVO enhanced by GPT-3.5 has a relatively low $\mathsf{Q}_{\mathsf{acc}}$ and has nearly 10% $\mathsf{Q}_{\mathsf{rej}}$ in critical error during the population-level crossover. Figure 5 also exhibits the similar result as format and critical check. The ratio of $\mathsf{Q}_{\mathsf{acc}}$ in moderate error of GPT-4 is higher than that obtained by GPT-3.5. It can be concluded that the quality of LLMs output is promising but strongly dependent on model capability. In phases that demand processing large data amount, such as population-level crossover and mutation, the LLM-based operator only achieves a maximum accuracy ($\mathsf{Q}_{\mathsf{acc}}$) of 92.3%. However, in phases requiring less data processing, it can maintain a minimum accuracy of 95%. There are also fewer errors for individual-level EVO than population-level EVO. We can conclude that the reliability of LLMs is mainly affected by the scale of the processed data.
Operation Dataset Format Error (GPT-4) Critical Error (GPT-4) Moderate Error (GPT-4)
$\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{acc}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$
Crossover Erods 99.3% 0.3% 0.4% 98.6% 1.4% 0.0% 50.0% 30.1% 19.3% 0.5%
Email 100.0% 0.0% 0.0% 99.3% 0.7% 0.0% 57.2% 24.0% 17.9% 0.8%
Astro 99.3% 0.7% 0.0% 98.3% 1.7% 0.0% 55.8% 26.2% 18.1% 0.8%
Mutation Erods 100.0% 0.0% 0.0% 99.3% 0.7% 0.0% 80.8% 14.6% 4.3% 0.1%
Email 98.0% 2.0% 0.0% 94.3% 4.3% 1.4% 82.1% 14.3% 3.2% 0.3%
Astro 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 83.3% 14.0% 2.7% 0.1%
TABLE XI: The validations of format and critical errors for the LLMs output during crossover and mutation.
IV-E Scalability analysis
The scalability of LLM-based operators is influenced the input data, which varies across different stages.
IV-E1 Initialization
As illustrated in Table VII, the effectiveness of candidate selection is significantly influenced by the size of the dataset. The input scale for LLMs during initialization (both candidate selection and sampling) increases linearly with the dataset size. For example, given a network of 10,000 nodes, we must input all their IDs with any specific metrics to LLMs for ranking, filtering, and sampling. It is not practical to input such large amounts of numerical data into LLMS and have them do such complex operations, not to mention the huge token overhead, meaning that LLMs may not be suitable for this data-intensive task.
Previous experiments indicate that LLMs are highly sensitive to the amount of input data. However, apart from the initialization phase, the input to LLMs is not the dataset itself but the population. Thus, the performance of LLM-based selection and reproduction relies more on hyperparameters than on the scale or type of dataset.
IV-E2 Selection
No matter what the dataset is, the input to LLM is always $\{(S_{1},f_{1}),...,(S_{k},f_{k})\}$. Thus, the relevant factor to LLM-based selection is the population size. A larger population size increases the amount of data LLMs need to process, but the common setting of this parameter is manageable to LLMs [77, 59]. Therefore, LLM-based selection has great applicability in this decision-making task given the promising result in Figure 3.
IV-E3 Crossover and Mutation
For these two phases, the input to LLMs is either $[\text{X}_{1},\text{X}_{2},...,\text{X}_{n}]$ (individual-level) or $\{[\text{X}_{1}^{(1)},\ldots,\text{X}_{n}^{(1)}],\ldots,[\text{X}_{1}^{(k)},\ldots,\text{X}_{n}^{(k)}]\}$ (population-level). As such, the input data amount, the main factor affecting LLMs’ performance, is dependent on the solution size and population size (only applicable to population-level optimization). The population size and solution size are empirical and will unnecessarily increase with the increase in the scale of datasets. Thus, we can conclude that the dataset scale and the LLM’s performance are not directly related, which is demonstrated in Table XI where the reliability of LLM-based EVO in the large dataset Astro is comparable to that in the two smaller networks. Evidence of LLMs sensitivity to hyperparameters can be found in Table XII, where the population size is set to $\mathcal{P}_{1}=30$ and $\mathcal{P}_{2}=10$, and the solution size is set to $\mathcal{S}_{1}=10$ and $\mathcal{S}_{2}=5$. Reducing either the population size or the size of a single solution helps to minimize errors, a trend consistent for both crossover and mutation phases.
One potential factor affecting performance as dataset size increases is the representation of element IDs in the solution. For example, given two datasets with element counts of up to $10^{3}$ and $10^{4}$, the maximum number of digits required to represent element IDs is 4 and 5, respectively. It means that the growth in element ID digits increases logarithmically rather than proportionally with dataset size. Therefore, we can conclude the scalability of LLM-based EVO in the reproduction phase is excellent, which is also demonstrated in Table XI, where there is no obvious difference between Astro and other smaller datasets regarding three types of check.
Netscience Format Error (GPT-4) Critical Error (GPT-4) Moderate Error (GPT-4)
$\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{app}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rep}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{acc}}}$ $\boldsymbol{\mathsf{Q}_{\mathsf{rej}}}$
C-$\boldsymbol{(\mathcal{P}_{1},\mathcal{S}_{1})}$ 96.7% 0.0% 3.3% 97.6% 2.4% 0.0% 51.8% 27.0% 20.4% 0.8%
C-$\boldsymbol{(\mathcal{P}_{1},\mathcal{S}_{2})}$ 100.0% 0.0% 0.0% 99.4% 0.6% 0.0% 73.3% 18.9% 7.6% 0.2%
C-$\boldsymbol{(\mathcal{P}_{2},\mathcal{S}_{1})}$ 99.3% 0.7% 0.0% 99.3% 0.7% 0.0% 93.2% 4.8% 1.5% 0.5%
M-$\boldsymbol{(\mathcal{P}_{1},\mathcal{S}_{1})}$ 100.0% 0.0% 0.0% 100.0% 0.0% 0.0% 77.7% 18.4% 3.8% 0.1%
M-$\boldsymbol{(\mathcal{P}_{1},\mathcal{S}_{2})}$ 100.0% 0.0% 0.0% 99.6% 0.4% 0.0% 92.9% 6.3% 0.8% 0.0%
M-$\boldsymbol{(\mathcal{P}_{2},\mathcal{S}_{1})}$ 100.0% 0.0% 0.0% 98.0% 2.0% 0.0% 95.7% 3.8% 0.5% 0.0%
TABLE XII: The validations for different errors for the LLMs output with different settings during crossover and mutation.
IV-F Validation of moderate error
We have conducted an investigation about what specific moderate errors the LLM-base EVO would encounter during crossover and mutation. Only the population-level optimization is tested as individual-level optimization rarely produces errors, as observed from Table VIII and X. The errors can be found in Tables 1 and 2 in the main text. Here we list them out for reference.
  • Error 1: The population after the operation remains the same as before.
  • Error 2: The size of the population fails to meet the requirement.
  • Error 3: The new solution does not preserve the original number of nodes in the original solution.
  • Error 4: The new solution contains duplicated elements.
  • Error 5: The new solution contains invalid nodes not found in candidate nodes. (only applicable to mutation)
These errors will be examined sequentially according to their index. When error 1 occurs, it will request a new operation from LLM. For the rest of errors, we will attach the corresponding message to the newly generated population and input them to LLM for a repaired population. During the repair process, the population with moderate error from the last check will pass through to the next phase if format error or critical error occurs. Figure 6 presents the results of moderate error in mutation.
As observed, the selection of LLM is one of the main factors of the reliability. GPT-4’s improvement over GPT-3.5 in managing mutation errors is evident across all error types except error 4, with a marked increase in the proportion of solutions that meet quality standards without the need for further modifications. On the other hand, GPT-3.5 is more likely to produce the unacceptable population with format or critical errors especially when trying to repair error 1 and error 3 while it rarely happens for GPT-4. It can also be found the frequency of error varies, for example, error 3 and error 5 occur less frequently than others. Regardless of the LLM, error 1 happens most frequently although it can be repaired by GPT-4, it implies that LLM sometimes will be overwhelmed with the data and does not work at all. The increased effectiveness in handling errors by GPT-4 suggests advancements in the model’s capability to generate more robust and reliable solutions, possibly due to improved understanding and processing of complex scenarios.
Figure 7 illustrates the moderate errors encountered during the crossover of two different-sized networks, Netscience and Astro. In contrast to mutation (Figure 6), the $\mathsf{Q}_{\mathsf{acc}}$ ratio for LLM-based crossover is lower, suggesting that LLMs struggle more with crossover, possibly due to the need to manage two populations compared to the single population in mutation. We found that both networks exhibit a similar pattern despite their significant size difference, suggesting that network size has little impact on the performance of LLM-based EVO regarding reproduction and highlighting its excellent scalability.
Figure 6: The encountered moderate errors during population-level LLM-based mutation. The tested network is Netscience $(|V|=379)$.
Figure 7: The encountered moderate errors during population-level LLM-based crossover. GPT-4 is used as the LLM for testing. Two networks of different sizes are tested: Astro $(V=14,845)$ and Netscience $(|V|=379)$.
IV-G Ablation study of repair mechanism
Figure 8 presents the ablation results of implementing a repair mechanism on the population-level LLM-based EVO across three datasets. We only test the moderate error as format error and critical error are so severe that the optimization cannot proceed as normal, and lead to consistently poor outcomes. For both datasets, the inclusion of a repair mechanism consistently outperforms the absence of one, as indicated by the higher average fitness achieved across generations. The observed stagnation in the scenarios without a repair mechanism indicates the impact of errors on optimization and suggests the importance of repair procedures.
Figure 8: The ablation result of repair mechanism.
IV-H Computational overhead analysis
Figure 9 compares the token costs between individual- and population-level LLM-based EVO in Netscience. Initial cost refers to the tokens required for a single operation, while total cost includes the additional tokens for repairs. The results show that population-level EVO incurs much lower costs than the individual-level EVO for both crossover and mutation in the entire process, no matter what LLM is used. It can be seen that the repair costs for the population-level approach are higher than those for the individual-based method. Nevertheless, both theoretically and practically, the total costs of the population-level method are lower. As LLMs continue to advance, the accuracy of LLM-based EVO will improve, leading to reduced repair costs, which further show the superiority of population-level optimization.
Figure 9: The computational cost of different optimization modes.
V Conclusion
To facilitate evolutionary optimization by LLMs, we have conducted a comprehensive evaluation of LLM-based EVO by establishing a set of criteria. Our assessment offers insights and analysis on the strengths and limitations of using LLMs as black-box EVO. Moreover, we have developed an original and novel population-level optimization approach using LLMs, and our proposed repair mechanism to mitigate uncertainties in LLM-generated outputs is also significant. This innovation enhances reliability and has the potential to be extended to numerous practical applications.